OpenStudy (maddieayeeee):

The Eagles won 3/5 of their regular season games and 4/5 of their tournament games. They played 15 regular seasons games and 5 tournament games. How many games did the Eagles win?

OpenStudy (wckeller):

15games * (3/5 won) = 9 games won 5tgames * (4/5 won) = 4 games won 9 + 4 = 13 games won in total.

OpenStudy (maddieayeeee):

Thank you so much!!

OpenStudy (wckeller):

Np